OpenStudy (anonymous):

Please can someone assist me??? Find all x in the interval (0, pi) that satisfy cot(x/2) > 1 + cot(x)

OpenStudy (curry):

You would need to use half angle formulas. Do you know them?

OpenStudy (anonymous):

the umm....sin(x/2) = SQRT (1-cosx)/2...?

OpenStudy (anonymous):

or the one of tan(x/2) = sin(x)/1+cos(x)?

OpenStudy (anonymous):

then for cot, the inverse is used yeah?